# A Survey of AI Agent Protocols

### v2

@@ Abstract
The rapid development of large language models (LLMs) has led to the widespread deployment of LLM agents across diverse industries, including customer service, content generation, data analysis, and even healthcare. However, as more LLM agents are deployed, a major issue has emerged: there is no standard way for these agents to communicate with external tools or data sources. This lack of standardized protocols makes it difficult for agents to work together or scale effectively, and it limits their ability to tackle complex, real-world tasks. A unified communication protocol for LLM agents could change this. It would allow agents and tools to interact more smoothly, encourage collaboration, and triggering the formation of collective intelligence. In this paper, we provide the first comprehensive analysis of existing agent protocols, proposing a systematic two-dimensional classification that differentiates context-oriented versus inter-agent protocols and general-purpose versus domain-specific protocols. Additionally, we conduct a comparative performance analysis of these protocols across key dimensions such as security, scalability, and latency. Finally, we explore the future landscape of agent protocols by identifying critical research directions and characteristics necessary for next-generation protocols. These characteristics include adaptability, privacy preservation, and group-based interaction, as well as trends toward layered architectures and collective intelligence infrastructures. We expect this work to serve as a practical reference for both researchers and engineers seeking to design, evaluate, or integrate robust communication infrastructures for intelligent agents.
@@ 2.2 Definition and Developments of Agent Protocols
ProtocolInter-AgentDomain-SpecificSystem-AgentAgent Protocol,LMOSHuman-ComputerLOKA, PXPRobot-AgentCrowdES, SPPsGeneral-PurposeANP, A2A, AITP, AComP, AConP, AgoraContext-OrientedDomain-Specificagents.jsonGeneral-PurposeMCP
Figure 3: Classification of various agent protocols from two dimensions, i.e., object orientation and application scenario. Please refer to Table 2 for more information.
@@ 3.1 Context-Oriented Protocols
Context- Oriented Genreal- Purpose MCP Anthropic (2024) Anthropic Connecting agents and resources RPC, OAuth Factual Standard
@@
AComP Al and Data (2025) IBM Multi agent system communication OpenAPI Drafting
@@ 3.2.1 General-Purpose Protocols
Several inter-agent protocols have already been proposed, including the Agent Network Protocol (ANP) (Chang, 2024), Google’s Agent2Agent Protocol (A2A) (Google, 2025), the Agent Interaction & Transaction Protocol (NEAR, 2025), the Agent Connect Protocol (AConP) (Cisco, 2025), and the Agent Communication Protocol (AComP) (Al and Data, 2025). Although all of them construct protocols focused on the interaction of agents, they vary in their problem domains, application scenarios, and implementation strategies. The following parts will discuss protocols above.
Table 3: Comparison of different inter-agent protocols. Development Stages assessed in Apr. 2025.
@@
AConP (Agent Connect Protocol) Standardize Interface to Invoke and Configure Agents Agents on Local Area Networks OpenAPI, JSON Drafting
AComP (Agent Communication Protocol) Standardize practical, valuable communication features Agents on Local Area Networks OpenAPI Drafting
@@
Agent Communication Protocol (AComP) (Al and Data, 2025)
It is a protocol designed to standardize how agents communicate, enabling automation, agent-to-agent collaboration, UI integration, and developer tooling. Rather than imposing strict specifications immediately, AComP emphasizes practical, useful features first, and will standardize features that demonstrate value, ensuring broader adoption and long-term compatibility. The motivation of AComP is that current agent systems often use diverse communication standards, causing complexity, integration difficulties, and vendor lock-in. To address these issues, AComP uniquely standardize interactions tailored specifically for agents which handle natural language inputs and depend on externally hosted models. AComP would like to simplify integration and promote effective collaboration across agent-based ecosystems by accommodating these agent-specific needs, but is still in the design.

### v3

@@ Abstract
The rapid development of large language models (LLMs) has led to the widespread deployment of LLM agents across diverse industries, including customer service, content generation, data analysis, and even healthcare. However, as more LLM agents are deployed, a major issue has emerged: there is no standard way for these agents to communicate with external tools or data sources. This lack of standardized protocols makes it difficult for agents to work together or scale effectively, and it limits their ability to tackle complex, real-world tasks. A unified communication protocol for LLM agents could change this. It would allow agents and tools to interact more smoothly, encourage collaboration, and triggering the formation of collective intelligence. In this paper, we provide the first comprehensive analysis of existing agent protocols, proposing a systematic two-dimensional classification that differentiates context-oriented versus inter-agent protocols and general-purpose versus domain-specific protocols. Additionally, we conduct a comparative performance analysis of these protocols across key dimensions such as security, scalability, and latency. Finally, we explore the future landscape of agent protocols by identifying critical research directions and characteristics necessary for next-generation protocols. These characteristics include adaptability, privacy preservation, and group-based interaction, as well as trends toward layered architectures and collective intelligence infrastructures. We expect this work to serve as a practical reference for both researchers and engineers seeking to design, evaluate, or integrate robust communication infrastructures for intelligent agents. Ongoing updates and the latest developments will be maintained at https://github.com/zoe-yyx/Awesome-AIAgent-Protocol.
@@ 2.2 Definition and Developments of Agent Protocols
ProtocolInter-AgentDomain-SpecificSystem-Agent(Sec. 3.2.2.3)Agent Protocol,LMOSRobot-Agent(Sec. 3.2.2.2)CrowdES, SPPsHuman-Computer(Sec. 3.2.2.1)LOKA, PXP, WAPGeneral-PurposeANP, A2A, AITP, AComP,AConP, Agora, CoralContext-OrientedDomain-Specificagents.jsonGeneral-PurposeMCP
Figure 3: Classification of various agent protocols from two dimensions, i.e., object orientation and application scenario. Please refer to Table 2 for more information. To distinguish between the two ACPs — the Agent Connect Protocol and the Agent Communication Protocol — we will refer to them as AconP and AcomP, respectively.
@@ 3.1 Context-Oriented Protocols
Context- Oriented General- Purpose MCP Anthropic (2024) Anthropic Connecting agents and resources RPC, OAuth Factual Standard
@@
AComP AI and Data (2025) IBM Multi agent system communication OpenAPI Drafting
@@
Coral citeagentcoralprotocol The Coral Community Multi agent system communication - Drafting
@@ 3.2.1 General-Purpose Protocols
Several inter-agent protocols have already been proposed, including the Agent Network Protocol (ANP) (Chang, 2024), Google’s Agent2Agent Protocol (A2A) (Google, 2025), the Agent Interaction & Transaction Protocol (NEAR, 2025), the Agent Connect Protocol (AConP) (Cisco, 2025), and the Agent Communication Protocol (AComP) (AI and Data, 2025). Although all of them construct protocols focused on the interaction of agents, they vary in their problem domains, application scenarios, and implementation strategies. The following parts will discuss protocols above.
Table 3: Comparison of different general-purpose inter-agent protocols. Development Stages assessed in Apr. 2025.
@@
AConP (Agent Connect Protocol) Standardize Interface to Invoke and Configure Agents Agent Communication and Value Exchange OpenAPI, JSON Drafting
AComP (Agent Communication Protocol) Standardize Practical, Valuable Communication Features Agent Communication and Value Exchange OpenAPI Drafting
Coral Decentralized Collaborative Infrastructure for AI Agents The Internet of Agents - Drafting
Agora Meta Protocol Communication Flexibility Inter Agent Collaboration Oxford Drafting
@@
Agent Communication Protocol (AComP) (AI and Data, 2025)
It is a universal protocol which allow agents to communicate and collaborate across different teams, frameworks, technologies and organizations, transforming the fragments landscape of today’s AI Agents into inter-connected team mates. Rather than imposing strict specifications immediately, AComP emphasizes practical, useful features first, and will standardize features that demonstrate value, ensuring broader adoption and long-term compatibility. It doesn’t manage workflows, deployments, or coordination between agents. Instead, AcomP enables orchestration across diverse agents by standardizing how they communicate. The currently lastest version holds four key features: REST-based communication (use HTTP patterns), no SDK required (but available), offline discovery and async-first, sync supported, enabling RESTful, streaming-compatible architecture, structured multipart messages (MIME) and token-based security.
The Coral Protocol (Georgio et al., 2025)
The Coral Protocol is an open and decentralized collaborative infrastructure designed to address the interoperability challenges among proprietary AI agents deployed by different organizations. It provides a unified framework for communication, coordination, trust, and payments among AI agents, enabling efficient collaboration across diverse domains and vendors. Key features of the protocol include a threaded communication model, the use of the Coral Server as a coordination layer, and MCP as a standardized agent communication protocol. Additionally, agents are defined by a structured identity, an assigned role, and scoped memory, which establish how agents interact with the system and with one another, enabling trust, modularity, and controlled information sharing. The protocol supports secure and efficient collaboration for agents built on any framework, positioning itself as a potentially practical protocol for integrating agents into software systems in the future.
@@ 3.2.2.1 Human–Agent Interaction Protocol
Web-Agent Protocol (OTA Tech AI, 2025)
The Web-Agent Protocol (WAP) defines a standardized interaction framework for coordinating user–agent–browser interactions through record-and-replay mechanisms. Designed to enhance the automation and auditability of web-based tasks, WAP decouples the recording of user interactions from their execution, enabling efficient reuse and precise task reproduction. The protocol operates in four stages: (1) user interactions are captured using the OTA-WAP Chrome Extension; (2) a backend service processes the raw event stream into structured action lists, supporting both exact and intelligent replay; (3) the action sequences can be embedded within contextual environments using the MCP (Message Context Protocol) server; and (4) these actions are executed in the browser via the WAP-Replay component. Key use cases include web automation, agent-based UI testing, and behavioral modeling for intelligent assistants. By providing a modular and interpretable way to integrate human inputs into web agents’ workflows, WAP complements existing intelligibility protocols and contributes to the development of responsive and accountable web-based agent systems.